The measure of an angle is 10 less than three times its complement. What is the measure of the two angles?
The equation I came up with was 3x-10=90. Solved for x and got 33.3. 90-33.3=56.7 but the answers are 65 and 25, not 33.3 and 56.7

Answer by mananth(16946)

let angle =x
complement = 90-x
3 times
=3(90-x)
...
x=3(90-x)-10
x=270-3x-10
4x=260
/4
x=65
the other angle =25